freddie is five times as old as elena, but in four years time, he will be 8 years more than thrice as old as her. Find the age of each now.
Answer by ikleyn(52803)

.
freddie is five times as old as elena, but in four years time, he will be 8 years more than thrice as old as her. Find the age of each now.
~~~~~~~~~~~~~~~~~~~~~~~~

F = 5E,           (1)
F+4 = 8+3(E+4).   (2)

It is your model.

From (1), substitute F = 5E into (2) replacing F. You will get

5E + 4 = 8 + 3E + 12.

Now simplify and solve.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".